# references.bib

### v1

@@ A.3 Experiments
The proposed method requires access to training data. Are there ways to relax this assumption? We explained in the paper that, to appropriately link navigation and the execution of a manipulation policy, a policy mobilization method needs to be policy-aware. This means that the policy mobilization method needs to have some information about the manipulation policy. Our proposed method assumes to have access to the initial observations of the training demonstrations. This is usually easy to obtain in open-source datasets and models. We mentioned in the Limitations section that there are ways to relax this assumption if the dataset is too large or not publically available. For example, you could sample a subset of demonstration episodes or rolling out the policy to collect some successful episodes. Note that the baselines also assume access to in-distribution data to different extents: BC w/ Nav requires collecting additional in-distribution data for policy learning and LeLaN is fine-tuned with in-distribution navigation data.
The proposed method has access to the 3D model of the test-time scene. How did you ensure a fair comparison to baselines that natively do not assume this information? To ensure a fair comparison, we provide the LeLaN and the VLFM baselines with ground-truth 3D models of the scene for collision detection. This detail is also mentioned in Section 3.3 of the main paper. The BC w/ Nav cannot take into account a 3D model of the test-time scene. Therefore, we do not provide it.
B Method Details
B.1 Hyperparameters
@@
B.2 Implementation Details
Our method takes as input one image at a time but can be trivially extended to a set of input images. This is useful when the robot is equipped with multiple cameras for example at the wrist or from an additional third-person viewpoint. To compute a predicted score $K(p)$ in this case, we evaluate the $K_{\text{id}}$ and $K_{\text{obj}}$ scores for each view, take the max, and then compose with the $K_{\text{col}}$ score to obtain the final score.
@@ B.2 Implementation Details
C Baseline Details
LeLaN. In our experiments, we assume that the robot initializes at a base position where it can navigate in a straight line to the target object without collision. Therefore, we used a pre-trained LeLaN [hiroselelan] checkpoint without the collision loss objective. We tried the pre-trained checkpoint in simulation and found it to perform poorly. To improve the robustness of the baseline, we fine-tune the pre-trained checkpoint with a randomized navigation dataset. For this dataset, we collected 300 demonstrations in each training layout and task, totaling $300\times 5\times 5=7,500$ navigation demonstrations. Each demo has a different, procedurally generated scene texture. In each demonstration, we retrieve the ground-truth target object position, add an offset to this position so we find the nearest robot pose in front of the object without collision, and add random noise to the position and rotation of this pose. As a result, we obtain a dataset of non-policy-aware navigation data where the robot initializes in training scenes and navigates to objects that the robot will encounter at test time. We fine-tune LeLaN on this generated dataset for 20 episodes before deploying it into each test-time environment. We only fine-tune a single LeLaN checkpoint to be used for all tasks.
VLFM. We integrated VLFM [yokoyama2024vlfm] into our framework with several key enhancements. Images and segmentation masks from RoboCasa [robocasa2024] were sent to VLFM, which generated high-level navigation commands such as TURN LEFT, TURN RIGHT, and STOP, based on its algorithm using pretrained feature extractors and object detectors to build semantic value maps for its exploration phase. Following exploration, we replaced VLFM’s pretrained navigation module (originally trained in Habitat [savva2019habitat]) with our own implementation of RRT, utilizing a point cloud representation of the scene for collision checking. Rather than navigating to the closest point on the target object, we improved accuracy by selecting the median point within the object’s point cloud. Additionally, we enhanced the performance of the pretrained MobileSAM [zhang2023faster] detector by incorporating RoboCasa’s ground truth masks, which significantly improved segmentation accuracy.
BC w/ Nav. In simulation, we collect 1,500 demonstrations per task, where each demonstration includes navigation and manipulation chained together. This is 3$\times$ more demos than the number of demos used for training the pure manipulation policy. Note that each demo also has a longer horizon than demos used to train a manipulation policy because of the additional navigation steps. In real-robot experiments, we collect the same number of episodes to train BC w/ Nav as our method. We use the same algorithm, hyperparameter setting, and number of epochs to train this baseline as the manipulation policy used for mobilization.
D Real Robot Experiment Details
@@
Rollout. We develop a multiprocessed pipeline for evaluating trained Diffusion Policies [chi2023diffusionpolicy] in real robot setups. An inference process repeatedly receives observations from the Basler camera and performs iterative denoising to predict a sequence of 16 actions at a time. The predicted actions and the corresponding observation timestamps are sent to the rollout process. At every step of policy execution, the rollout process receives an action sequence from the inference process and then skips a few actions so that the predicted actions sent to the robot are aligned in time with what the policy expects the robot to execute. During evaluation, we use the same DDPM noise scheduler as the training procedure.
@@
Details of the Humans baseline. For each task, we ask 10 unique users to drive the mobile robot to where they believe is the optimal starting base pose to execute the task. All users have technical backgrounds in robotics but are not aware of how the policies are trained and what robot positions are best for policy execution. We then initialize the robot at each of these human-designated navigation targets, and then evaluate the manipulation policy from there. This allows us to test whether our method can even outperform navigation targets generated via human intuition.
@@
Measuring task success. For each task and method, we track the incidence rates of certain performance milestones (i.e., navigation success, grasp success, full success), and also of task ending-collisions (i.e., fatal failures). We consider a navigation attempt successful if the robot drives to a pose that not only faces the target object but also is within a distance of 50 cm from the ideal target pose. A grasp success is defined to be when the robot stably grasps the object of interest (i.e. the chip bag, the paper towels, or the tray of tomatoes), and full success is counted only if the entire task is completed.
@@
Detailed analysis of results. We find that for all three manipulation tasks, our method significantly outperforms the two baselines. The end-to-end BC w/ Nav baseline performs decently well in the Pour Tomatoes task, where the robot can navigate to a reasonable position $80\%$ of the time, ultimately achieving a full success rate of $70\%$ with only $20\%$ of the runs ending in collision. In the Bimanual Grab Paper Towels and Pick Chips tasks however, it sees an uptick in fatal failures at $70\%$ and $40\%$ respectively. In our experiments, BC w/ Nav only achieves a $30\%$ navigation success rate for both of these two tasks, which even further degrades into a $0\%$ full success rate for the Pick Chips task.
@@
The human baseline achieves a very high navigation success of at least $80\%$ for all tasks. However, the success rate of the entire task is very low, with a maximum score of $40\%$ in Bimanual Grab Paper Towels. In the Pick Chips task, $20\%$ of the rollouts even end in arm-collisions with the shelf. This result yields an interesting insight—although humans are adept at intuitively navigating the robot into very reasonable starting poses, they are ultimately unaware of what the policy training data looks like. Therefore, the human-provided navigation poses still generally fail to achieve task success. Our method is able to pick up on the subtleties of these distribution shifts. In real experiments, our method achieves successful final navigation poses $100\%$ of the time for all three tasks, and achieves full success rates of $70\%$, $100\%$, and $80\%$ for the Pick Chips, Bimanual Grab Paper Towels, and Pour Tomatoes tasks respectively. Furthermore, unlike any of the evaluated baselines, the rollouts using our method are completely collision-free.
E Additional Visualizations

### v2

@@ A.3 Experiments
The proposed method requires access to training data. Are there ways to relax this assumption? We explained in the paper that, to appropriately link navigation and the execution of a manipulation policy, a policy mobilization method needs to be policy-aware. This means that the policy mobilization method needs to have some information about the manipulation policy. Our proposed method assumes access to the initial observations of the training demonstrations. This is usually easy to obtain in open-source datasets and models. We mentioned in the Limitations section that there are ways to relax this assumption if the dataset is too large or not publicly available. For example, you could sample a subset of demonstration episodes or rolling out the policy to collect some successful episodes. Note that the baselines also assume access to in-distribution data to different extents: BC w/ Nav requires collecting additional in-distribution data for policy learning, and LeLaN is fine-tuned with in-distribution navigation data.
The proposed method has access to the 3D model of the test-time scene. How did you ensure a fair comparison to baselines that natively do not assume this information? To ensure a fair comparison, we provide the LeLaN and the VLFM baselines with ground-truth 3D models of the scene for collision detection. This detail is also mentioned in Section 6 of the main paper. The BC w/ Nav cannot take into account a 3D model of the test-time scene. Therefore, we do not provide it.
B Additional Ablation Results
Ours w/o $K_{\text{id}}$ w/o $K_{\text{col}}$ w/o $K_{\text{obj}}$
Mean Success Rate 0.422 0.006 0.408 0.344
Table 2: Ablations on hybrid score function. We compute the average success rate of each method across all 5 tasks, with a total of 750 evaluation episodes for each number shown in the paper.
To study the individual importance of each element in the hybrid score function, we add an ablation study where we compare our method with variants of our method without one of the three hybrid score components. In Table 2, we show that each component of the hybrid score function contributes to the final performance of our proposed method.
C Method Details
C.1 Hyperparameters
@@
C.2 Implementation Details
Our method takes as input one image at a time, but can be trivially extended to a set of input images. This is useful when the robot is equipped with multiple cameras for example at the wrist or from an additional third-person viewpoint. To compute a predicted score $K(p)$ in this case, we evaluate the $K_{\text{id}}$ and $K_{\text{obj}}$ scores for each view, take the max, and then compose with the $K_{\text{col}}$ score to obtain the final score.
@@ C.2 Implementation Details
D Baseline Details
LeLaN. In our experiments, we assume that the robot initializes at a base position where it can navigate in a straight line to the target object without collision. Therefore, we used a pre-trained LeLaN [17] checkpoint without the collision loss objective. We tried the pre-trained checkpoint in simulation and found it to perform poorly. To improve the robustness of the baseline, we fine-tune the pre-trained checkpoint with a randomized navigation dataset. For this dataset, we collected 300 demonstrations in each training layout and task, totaling $300\times 5\times 5=7,500$ navigation demonstrations. Each demo has a different, procedurally generated scene texture. In each demonstration, we retrieve the ground-truth target object position, add an offset to this position so we find the nearest robot pose in front of the object without collision, and add random noise to the position and rotation of this pose. As a result, we obtain a dataset of non-policy-aware navigation data where the robot initializes in training scenes and navigates to objects that the robot will encounter at test time. We fine-tune LeLaN on this generated dataset for 20 episodes before deploying it into each test-time environment. We only fine-tune a single LeLaN checkpoint to be used for all tasks.
VLFM. We integrated VLFM [18] into our framework with several key enhancements. Images and segmentation masks from RoboCasa [20] were sent to VLFM, which generated high-level navigation commands such as TURN LEFT, TURN RIGHT, and STOP, based on its algorithm using pretrained feature extractors and object detectors to build semantic value maps for its exploration phase. Following exploration, we replaced VLFM’s pretrained navigation module (originally trained in Habitat [93]) with our own implementation of RRT, utilizing a point cloud representation of the scene for collision checking. Rather than navigating to the closest point on the target object, we improved accuracy by selecting the median point within the object’s point cloud. Additionally, we enhanced the performance of the pretrained MobileSAM [94] detector by incorporating RoboCasa’s ground truth masks, which significantly improved segmentation accuracy.
[…]
E Real Robot Experiment Details
@@
Rollout. We develop a multiprocessed pipeline for evaluating trained Diffusion Policies [1] in real robot setups. An inference process repeatedly receives observations from the Basler camera and performs iterative denoising to predict a sequence of 16 actions at a time. The predicted actions and the corresponding observation timestamps are sent to the rollout process. At every step of policy execution, the rollout process receives an action sequence from the inference process and then skips a few actions so that the predicted actions sent to the robot are aligned in time with what the policy expects the robot to execute. During evaluation, we use the same DDPM noise scheduler as the training procedure.
@@
Details of the Humans baseline. For each task, we ask 10 unique users to drive the mobile robot to where they believe is the optimal starting base pose to execute the task. All users have technical backgrounds in robotics, but are not aware of how the policies are trained and what robot positions are best for policy execution. We then initialize the robot at each of these human-designated navigation targets, and then evaluate the manipulation policy from there. This allows us to test whether our method can even outperform navigation targets generated via human intuition.
@@
Measuring task success. For each task and method, we track the incidence rates of certain performance milestones (i.e., navigation success, grasp success, full success), and also of task-ending collisions (i.e., fatal failures). We consider a navigation attempt successful if the robot drives to a pose that not only faces the target object but also is within a distance of 50 cm from the ideal target pose. A grasp success is defined as when the robot stably grasps the object of interest (i.e. the chip bag, the paper towels, or the tray of tomatoes), and full success is counted only if the entire task is completed.
@@
Detailed analysis of results. We find that for all three manipulation tasks, our method significantly outperforms the two baselines. The end-to-end BC w/ Nav baseline performs decently well in the Pour Tomatoes task, where the robot can navigate to a reasonable position $80\%$ of the time, ultimately achieving a full success rate of $70\%$ with only $20\%$ of the runs ending in collision. In the Bimanual Grab Paper Towels and Pick Chips tasks, however, it sees an uptick in fatal failures at $70\%$ and $40\%$ respectively. In our experiments, BC w/ Nav only achieves a $30\%$ navigation success rate for both of these two tasks, which even further degrades into a $0\%$ full success rate for the Pick Chips task.
@@
The human baseline achieves a very high navigation success of at least $80\%$ for all tasks. However, the success rate of the entire task is very low, with a maximum score of $40\%$ in Bimanual Grab Paper Towels. In the Pick Chips task, $20\%$ of the rollouts even end in arm-collisions with the shelf. This result yields an interesting insight—although humans are adept at intuitively navigating the robot into very reasonable starting poses, they are ultimately unaware of what the policy training data looks like. Therefore, the human-provided navigation poses still generally fail to achieve task success. Our method is able to pick up on the subtleties of these distribution shifts. In real experiments, our method achieves successful final navigation poses $100\%$ of the time for all three tasks, and achieves full success rates of $70\%$, $100\%$, and $80\%$ for the Pick Chips, Bimanual Grab Paper Towels, and Pour Tomatoes tasks, respectively. Furthermore, unlike any of the evaluated baselines, the rollouts using our method are completely collision-free.
F Additional Visualizations